ELASTICITY

IN-CLASS WORKSHEET 2

This question examines the market for mangos. You will use a demand function to construct the demand schedule, calculate the price elasticity of demand at different points along a linear demand curve, and identify the likely effects of price changes on total revenue.

Below, you are provided with the demand function for mangos. If you plug any price into the formula for the demand function, you get the quantity demanded at that price.

Q = 150 – 25P

Task 1: Use the table below to find the quantity of mangos demanded at each price.

Price

Quantity of Mangos Demanded

$1

2

3

4

5

6

Task 2: Calculate the price elasticity of demand when the price falls from $5 to $4.

Task 3: When the price of a mango falls from $5 to $4, does total revenue fall or rise? How do you know?

Task 4: When the price of a mango falls from $3 to $2, does total revenue fall or rise?
Answer #1

Price

Quantity of Mangoes Demanded

$1

150-25(1)=125

2

150-25(2)=100

3

150-25(3)=75

4

150-25(4)=50

5

150-25(5)=25

6

150-25(6)=0

Task 2: Quantity demanded at $5 (Q1) = 25

Quantity demanded at $4 (Q2) = 50

(02+8) Price 2 elasticity demand – - P 2 25-50 (25+50 2 - 12 = 3 4-5 4+5 2

Task 3: When price falls from $5 to $4, the quantity demanded rises from 25 to 50. This means that the total revenue rises (from the table)

Task 4: When price falls from $4 to $3, the quantity demanded rises from 50 to 75. This also means that the total revenue rises. The increase in demand stays constant after each dollar price fall. This means that the slope of the demand curve is constant.